OpenStudy (anonymous):

If (a, 1996) and (1996, b) are different points on the graph of the line y=3/4x+5, what is the value of 1996-a/1996-b ?

OpenStudy (anonymous):

1996=3/4a+5 solve a b=3/4(1996)+5 solve b